What is the slope of (2,3)and (9,-5)?
Answer by lwsshak3(11628)

What is the slope of (2,3)and (9,-5)?
slope=change in y/change in x=rise/run=∆y/∆x
For given points:
slope=(-5-3)/(9-2)=-8/7